OpenStudy (anonymous):

Find A so that 3log(base6)625 - 3log(base 6)5 = log(base 6)A

OpenStudy (anonymous):

3log(base6)(625/5)=log(base6)A log(base6)(625/5)^3=log(base6)A A=(625/5)^3

OpenStudy (anonymous):

Thanks